Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name Ole Hansen
Log entry time 10:01:05 on March30,2011
Entry number 353719
keyword=Magnet status
All cryo levels are back, except for R-Q1. It will take a while for the vacuum to get better before the magnet can be cooled again.
The vacuum problem in R-Q1 apparently caused a ripple effect, tripping several of the other quads. Current, L-Q1, L-Q3, and all RHRS quads are down and can't be reset from upstairs.